two consecutive integers with half of the smaller equual to 4 more than one third of the larger
Answer by mananth(16949)

let one integer be x
the other will be x+1
1/2 x = 1/3 *(x+1) +4
x/2- 4 = 1/3* (x+1)
(x/2 -4) =1/3(x+1)

x/2 -4 = x/3 + 1/3
x/2 - x/3 = 4+1/3
3x-2x / 6 = 13/3
3x=78
x=26
the other number is 27